# About the Burton-Miller factor in the low frequency region

### v1

@@ 3.1.1 Unit Sphere
Table 1: Rounded theoretical and numerically determined (up to three digits) critical frequencies for the sound hard unit sphere with $N=1280$ and $N=5120$ elements. The speed of sound was set to $c=340$ m/s. The numerically critical frequencies were determined by looking at the condition number of the BIE.
@@
In Fig. 4, the condition numbers at $f=170.656$, $200,244.083$, and $343.315$ Hz are depicted as function of the Burton-Miller factor $\eta$ for $N=1280$ elements. Especially for low frequencies, $\eta=\frac{{\mathrm{i}}}{k}$ (denoted by the dotted line) is not the optimal choice. This may be because a) of the geometric error introduced by the mesh, which means that the solution of the BEM is different from the analytic solution, b) the frequencies investigated were not the exact critical frequencies and c) that the wavenumber dependent parts of the derivatives of the Green’s function were too small compared to the $O(\frac{1}{r})$ and $O(\frac{1}{r^{2}})$ factors. In general, most papers dealing with finding a quasi optimal Burton-Miller factor look at the unit sphere and the analytic solution and eigenvalues for the non-discretized sphere, respectively. As relatively few elements are used for the BEM, there is already some difference in geometry to the “optimal” case. Fig. 4b, where a non-critical frequency was investigated, also shows that the non Burton-Miller version $\eta=0$ is better conditioned for the sound hard sphere.
@@
In Fig. 5b, the residuum of the Burton-Miller solution inserted into the original BIE is depicted. Although the Burton-Miller system is solved with LAPACK to an accuracy up to $10^{-16}$, the solution of the BIE itself is different from the Burton-Miller solution, which is mirrored in the residuum. The different settings for $\eta$ influence this residuum to a large degree, again, with the additional iteration steps this residuum can be kept relatively small.
@@
3.2 Sphere with $r=2$, $N=5120$
As the BM-factor is often derived by investigating the unit sphere as the scatterer (cf. [1, 2, 11, 19, 21]), the influence of the size of the scatterer will be investigated in this subsection. In the following, condition numbers, residuum, and relative errors for the “classical” BM-factor $\eta=\frac{{\mathrm{i}}}{k}$, for $\eta=\frac{{\mathrm{i}}}{2k}$, the factors proposed by [5] and [9], and the data for the iteration approach will be investigated for the example of a sound hard sphere with radius $r=2$.
Figure 8: a) Condition numbers and b) mean relative error for different choices of the BM factor for the sphere with radius $r=2$. The sphere was discretized using $N=5120$ elements.
It is clear that for $r=2$ the number of critical frequencies in the observed frequency range grows as they now depend on the roots of $j_{\ell}(2k),\ell=0,1,\dots$. Thus, the condition number of the BEM without Burton-Miller grows earlier illustrating the need for having the regularization provided by the BM-method.
For $r=2$ and $k>\pi$ the condition number for the non-Burton-Miller version is always higher than the number for the versions with Burton-Miller for all investigated choices of $\eta$ (see for example Fig. 8). The influence of the critical frequencies on the BIE becomes dominant for $k>\pi$, and if the BM-factor is interpreted as a weighting of BIE and dBIE, it is reasonable that the condition number for $\eta=\frac{{\mathrm{i}}}{2k}$ is bigger than for $\eta=\frac{{\mathrm{i}}}{k}$.
@@ 3.3 Sound Soft Sphere
Also, it is not trivial to construct one single and “cheap” iteration step to increase the accuracy of the BM solution anymore. Theoretically, it is possible to use the dBIE to construct the iteration
${\bf v}_{n+1}=2\left({\bf v}_{\text{inc}}-{\bf H}^{\prime}{\bf v}_{n}\right),$
however, since the stability of the dBIE and its solution becomes bad for $f\rightarrow 0$, this step does not result in any error reduction. It is, on the other hand, possible to use one or two steps of a Richardson Iteration applied to the BIE
@@
but a single step does not yield a substantial gain in accuracy as Fig. 9 illustrates.
Hinzufügen Kommentar zu ill posedness von diesem Problem
@@ 4 Discussion
Regularization methods like the Burton-Miller method play an important role for the BEM in acoustics. In general, the frequency dependent coupling factor $\eta=\frac{{\mathrm{i}}}{k}$ is a good choice for coupling BIE and dBIE to regularize the solution of the BIE for all frequencies. However, the numerical experiments made in the previous sections showed that especially in the low frequency region, where the critical frequencies are sparse and the numerical instabilities of the BIE are concentrated around those frequencies, a different choice for the BM factor can have a positive effect on the condition number and in turn on the calculation error. Originally, the coupling factor $\eta=\frac{{\mathrm{i}}}{k}$ has been derived by investigating the condition number of scattering problems involving the unit sphere, but there is always the difference between numerical approximation using, e.g., the BEM, and the analytic representation of the solution using spherical harmonics. This difference can be seen, for example, when comparing the critical frequencies of the analytic approach with the frequencies with high condition numbers for the numerical approximation.
When looking at Fig. 3, it becomes clear that apart from the critical frequencies, the condition numbers in the low frequency region for the BIE is smaller than for the Burton-Miller formulation.
@@
Using a relatively small BM factor in the low frequency regions as, e.g., proposed by [9], or for larger objects the factor dependent on the cross-section of the scatterer as proposed by [5] has a positive effect on the condition number and the accuracy of the solution. The factors proposed by [5, 9] have the slight drawback, that their definition changes at a certain frequency, and it is not clear, how this frequency is motivated. In [9] this frequency was defined by $\max(|k|^{2},50)$, which works excellent for the unit sphere (cf. Fig. 3), but leads to higher than necessary condition numbers for a bigger sphere (see for example Fig. 8).
As an alternative to using different BM factors for different frequency regions, it was also investigated how one additional step of a modified Richardson iteration for the BIE using the BM-solution with $\eta=\frac{{\mathrm{i}}}{k}$ as a starting value increases the accuracy of the BM solution. This approach is computationally relatively cheap, as it involves only one additional matrix-vector multiplication plus two vector additions. Of course, one may argue, that the additional iteration step may be unnecessary for higher frequencies, and that it would be hard to define beforehand when this additional iteration should be made or not. On the other hand, it would be easy to compare the solution vector before the additional iteration step and after. If the difference is below a certain tolerance, one could cease to use the additional iteration for frequencies higher than the current one. However, this additional iteration step only yields good errors for sound hard boundary conditions.

### v2

@@ 4.1.1 Sphere with radius $R=1$ m
Table 1: Rounded theoretical and numerically determined (up to three digits) critical frequencies for the sound hard unit sphere with $N=1280$ and $N=5120$ elements. The numerically critical frequencies were determined by looking at the condition number of the BIE.
@@
In Fig. 4, the condition numbers at $f=170.656$, $200,244.083$, and $343.315$ Hz are depicted as function of the Burton-Miller factor $\eta$ for $N=1280$ elements. Especially for low frequencies, $\eta=\frac{{\mathrm{i}}}{k}$ (denoted by the dotted line) does not lead to the smallest condition number. This may be because a) of the geometric error introduced by the mesh, which means that the solution of the BEM is different from the analytic solution, b) the frequencies investigated were not the exact critical frequencies and c) that the wavenumber dependent parts of the derivatives of the Green’s function were too small compared to the $\mathcal{O}(\frac{1}{r})$ and $\mathcal{O}(\frac{1}{r^{2}})$ factors. In general, most papers dealing with finding a quasi optimal Burton-Miller factor look at the unit sphere and the analytic solution and eigenvalues for the non-discretized sphere, respectively. As relatively few elements are used for the BEM, there is already some difference in geometry to the “optimal” case. Fig. 4b, where a non-critical frequency was investigated, also shows that the non Burton-Miller version $\eta=0$ is better conditioned for the sound hard sphere.
@@
On a sidenote, the fact that a solution $\boldsymbol{\phi}_{\text{BM}}$ solves the discretized combined boundary integral equation with an relative residuum of about $10^{-16}$, does not necessarily mean, that this solution also solves the original discretized BIE with a similar accuracy. Fig. 5b illustrates this behavior of the Burton Miller solution that was calculated using a direct solver from LAPACK. For $f=300$ Hz, for example, the mean residuum of the Burton Miller solution $\boldsymbol{\phi}_{\text{BM}}$ inserted in the discretized BIE
${\boldsymbol{\varepsilon}}_{\text{BIE}}=\left(\frac{\bf I}{2}-{\bf H}\right)\boldsymbol{\phi}_{\text{BM}}-\boldsymbol{\phi}_{\text{inc}}$
is on average about $3\cdot 10^{-6}$, a similar range can be observed if $\boldsymbol{\phi}_{\text{BM}}$ is inserted in the dBIE:
${\boldsymbol{\varepsilon}}_{\text{dBIE}}=-{\bf E}\boldsymbol{\phi}_{\text{BM}}-{\bf v}_{\text{inc}}.$
Only the combination ${\boldsymbol{\varepsilon}}_{\text{BIE}}+\eta{\boldsymbol{\varepsilon}}_{\text{dBIE}}\approx 10^{-16}$ is in the range of the machine precision. Although not really relevant for the real error of the solution compared to the analytic solution, it is, at least in the opinion of the author, an interesting behavior of the Burton Miller method.
@@
4.2 Sphere with $R=2$ m, $N=5120$
As the BM-factor is often derived by investigating the unit sphere as the scatterer (cf. [2, 3, 12, 22, 24]), the influence of the size of the scatterer will be investigated in this subsection. In the following, condition numbers, residuum, and relative errors for the “classical” BM-factor $\eta=\frac{{\mathrm{i}}}{k}$, for $\eta=\frac{{\mathrm{i}}}{2k}$, the factors proposed by [6] and [10], and the data for the iteration approach will be investigated for the example of a sound hard sphere with radius $R=2$ m.
Figure 8: a) Condition numbers and b) mean relative error for different choices of the BM factor for the sphere with radius $R=2$ m. The sphere was discretized using $N=5120$ elements.
It is clear that for $R=2$ m the number of critical frequencies in the observed frequency range grows as they now depend on the roots of $j_{\ell}(2k),\ell=0,1,\dots$. Thus, the condition number of the BEM without Burton-Miller grows earlier illustrating the need for having the regularization provided by the BM-method.
For $R=2$ m and $k>\pi$ the condition number for the non-Burton-Miller version is always higher than the number for the versions with Burton-Miller for all investigated choices of $\eta$ (see for example Fig. 8). The influence of the critical frequencies on the BIE becomes dominant for $k>\pi$, and if the BM-factor is interpreted as a weighting of BIE and dBIE, it is reasonable that the condition number for $\eta=\frac{{\mathrm{i}}}{2k}$ is bigger than for $\eta=\frac{{\mathrm{i}}}{k}$.
@@ 4.3 Sound Soft Sphere
The approach proposed by Burton and Miller can be “readily extended to allow for complex values of $k$ and more general boundary conditions” [7]. It was also shown in [5, 14] that the non-uniqueness problem in connection with Dirichlet boundary conditions can be avoided by using a suitable combination of single- and double-layer potentials.
@@
Also, it is not trivial to construct one single and “cheap” iteration step to increase the accuracy of the BM solution anymore. As already mentioned in Section 3 an iteration based on the dBIE is not recommended because of stability problems. However, it is possible to use one or two steps of a Richardson Iteration applied to the BIE
@@
but for the reason mentioned above the Burton Miller solution is not a good starting point for small $k$ and one single Richardson Iteration step does not yield a substantial gain in accuracy as Fig. 9 illustrates.
@@ 5 Discussion
Regularization methods like the Burton-Miller method play an important role for the BEM in acoustics. In general, the frequency dependent coupling factor $\eta=\frac{{\mathrm{i}}}{k}$ is a good choice for coupling BIE and dBIE to regularize the solution of the BIE for most frequencies. However, the numerical experiments made in the previous sections showed that especially in the low frequency region, where the critical frequencies are sparse and the numerical instabilities of the BIE are concentrated around those frequencies, a different choice for the BM factor can have a positive effect on the condition number and in turn on the calculation error. Originally, the coupling factor $\eta=\frac{{\mathrm{i}}}{k}$ has been derived by investigating the condition number of scattering problems involving the unit sphere for relatively high frequencies, but there is always the difference between numerical approximation using, e.g., the BEM, and the analytic representation of the solution using spherical harmonics. This difference can be seen, for example, when comparing the critical frequencies of the analytic approach with the frequencies with high condition numbers for the numerical approximation.
When looking at Fig. 3, it becomes clear that apart from the critical frequencies, the condition numbers in the low frequency region for the BIE is smaller than for the Burton-Miller formulation. From the eigenvalue analysis in Section 2 it becomes clear that for the problem on the sphere the dBIE and thus the Burton-Miller system becomes unstable if $k\rightarrow 0$, which explains the smaller condition number for the version without using the Burton-Miller method.
@@
Using a BM factor in the low frequency region that is not dependent on $\frac{1}{k}$ as, e.g., proposed by [10] or for larger objects the factor dependent on the cross-section of the scatterer as proposed by [6] has a positive effect on the condition number and the accuracy of the solution as also illustrated by the eigenvalue analysis for a sphere in Section 2. The factors proposed by [6, 10] have the slight drawback, that their definition changes at a certain frequency, and it is not clear, how this frequency is motivated. In [10] this frequency was defined by $\max(|k|^{2},50)$, which works excellent for the sphere of radius $R=1$ m (cf. Fig. 3), but leads to higher than necessary condition numbers for a bigger sphere (see for example Fig. 8). Nevertheless, as a one-glove-fits-it-all BM factor the factor proposed in [10] is a very good choice. Compared to the usual choice $\eta=\frac{{\mathrm{i}}}{k}$ it has the advantage that the systems for both sound hard and sound soft problems stay stable for small $k$ (see Eq. (7)).
As an alternative to using different BM factors for different frequency regions, it was also investigated how one additional step of a modified Richardson iteration for the BIE using the BM-solution with $\eta=\frac{{\mathrm{i}}}{k}$ as a starting value increases the accuracy of the BM solution. This approach is computationally relatively cheap, as it involves only one additional matrix-vector multiplication plus two vector additions. Of course, one may argue, that the additional iteration step may be unnecessary for higher frequencies, and that it would be hard to define beforehand when this additional iteration should be made or not. On the other hand, it would be easy to compare the solution vector before the additional iteration step and after. If the difference is below a certain tolerance, one could cease to use the additional iteration for frequencies higher than the current one. However, this additional iteration step only yields low errors for sound hard boundary conditions.
6 Acknowledgements
The author would like to thank the reviewers for their constructive remarks, which helped to improve the paper a lot.